OpenStudy (anonymous):

Use U-Substitution to evaluate the Integral:

OpenStudy (anonymous):

u=(x^{4} + 3x^{2} +5)

OpenStudy (anonymous):

\[\int\limits_{?}^{?} (4x^{3} +6x) \cos (x^{4} + 3x^{2} + 5 ) dx\]

OpenStudy (anonymous):

I have to integrate that above this post with the U given in the first post. Please confirm my answer if you can!!

OpenStudy (anonymous):

Actually I couldn't get an answer - please helpz!!!

OpenStudy (anonymous):

= sin ( u ) = sin ( x^4 + 3x^3 + 5) + C

OpenStudy (anonymous):

sorry the second x should be squared ( 3x^2)

OpenStudy (anonymous):

because u = x^4 + 3x^3 + 5 ==> du/dx = 4x^3 + 6x solve for dx and plug the equation for dx into the integral

OpenStudy (anonymous):

if you write that out you will notice you have 4x^3 + 6x in the numerator and denominator and that equals 1. but you are still multiplying by du. because you had to solve for dx

OpenStudy (anonymous):

you get \[\int\limits_{?}^{?}\cos u du\] = sin u then just plug u back into the sing

OpenStudy (anonymous):

taking a step back when you solve for dx you should get dx = du / (4x^3 + 6x) plug that into the integral. you are integrating du with respect to u

OpenStudy (anonymous):

\[\sin(x^{4} + 3x^{2} + 5) + c \]

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

awesome!! thank you so much!!

OpenStudy (anonymous):

yeahp!